OpenStudy (anonymous):

medal reward and fan! Evaluate –x2 + 2x – 4 with x = –2.

OpenStudy (anonymous):

. -(-2^2) + 2(-2) – 4= -12

OpenStudy (anonymous):

Evaluate 2x2 + 6x + 30 with x = –4. 86 54 38 24 7. Evaluate 6x2 – 2x + 7 with x = –10. 627 11 -627 -11 8. Evaluate –4x2 – 9x – 8 with x = –3.

OpenStudy (anonymous):

don't know...